(Density property) if a, b ∈ R with a < b, then there exists a rational number r ∈ Q such that a < r < b.


1
Expert's answer
2021-01-18T15:33:51-0500

Here as "b-a>0" , therefore using the Archimedean property we have, "n\in \Z_{>0}" such that

"n(b-a)>1" .


Now, let "m=[na]" (g.i.f). Then "m\leq na<m+1" . Also we have "m+1<ny" . (To get this you can assume "m+1\geq ny" and arrive at a contradiction)


Thus we have,


"m\leq na<m+1<nb"

Dividing all sides by "n",



"\frac{m}{n}\leq a < \frac{m+1}{n}<b"

Therefore as "m,n\in\Z" , we have "r=\frac{m+1}{n}\in \mathbb{Q}" as the rational number between "a" and "b" .